OpenStudy (anonymous):

Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

@Vocaloid @UsukiDoll

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

please help @Loser66

OpenStudy (loser66):

Can you take a snapshot?

OpenStudy (anonymous):

of equations?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

does that help?

OpenStudy (loser66):

yes, I can see it. Let me think

OpenStudy (anonymous):

I have no clue what im doing

OpenStudy (anonymous):

@Mehek14 can you help

OpenStudy (anonymous):

can anyone help because i am lost and i think you have a much better idea whats going on here than i do

OpenStudy (anonymous):

@Vocaloid @UsukiDoll

OpenStudy (loser66):

Maybe, I overthink of it. May be, it is just f(x) = x +2 g(x) = x -2 since then f(g(x) = x and g(f(x)) = x also and then f , g are inverse of each other.!! ha!!

OpenStudy (loser66):

if it is that simple, then a =2, d = 2 , c = 1 , b =1.

OpenStudy (loser66):

@freckles Am I underestimate the problem?

OpenStudy (freckles):

\[f(x)=\frac{x+a}{b} \\ b f(x)=x+a \\ b f(x)-a=x \\ f^{-1}(x)=bx-a \\ \text{ is suppose to be equal to } g(x) \\ bx-a=cx-d \\ \implies b=c \text{ and } a=d\] the possibilities are infinite

OpenStudy (freckles):

so loser's values will work